OpenStudy (anonymous):

The following is a geometric sequence. 1, -2, 4, -8, 16,

OpenStudy (triciaal):

multiply by -2

OpenStudy (jmark):

For the given sequence r = T2/T1 = -2/1 = -2 = T3/T2 = 4/-2 = -2 = T4/T3 = -8/4 = -2 The difference between the term is same. The given sequence is in geometric progression. ICSE syllabus for class 6 - http://icse.edurite.com/icse-syllabus/icse-syllabus-for-class-6.html